SPECTROSCOPIC INVESTIGATION OF ELECTRON-INDUCED PROTON TRANSFER IN THE FORMIC ACID DIMER, (HCOOH)_{2}

\underline{HELEN K. GERARDI}, CHRIS M. LEAVITT, ANDREW F. DEBLASE, AND MARK A. JOHN-SON, Yale University, Department of Chemistry, New Haven, CT.

We have isolated the stable form of the formic acid dimer anion (HCOOH)_{2}^{−}, a model for electron-induced proton transfer between nucleic acid base-pairs, in the gas phase. The vibrational signatures of this species and its various isotopomers were investigated using Ar predissociation and photodetachment spectroscopies in the 600-3800 cm^{−1} range. We relate the experimental infrared transitions of the anion to those predicted for its calculated lowest energy structure in order to determine if a proton transfer event does in fact occur upon excess electron attachment to this simple hydrogen-bonded dimer. Additionally, we determined its vertical detachment energy (VDE), 1.8 eV, using velocity-map photoelectron imaging.